OpenStudy (anonymous):

could i find two missing angle measures if i know some of the side length of a right triangle? (can i have an example ?? )

OpenStudy (anonymous):

xD

OpenStudy (anonymous):

?

OpenStudy (anonymous):

please help me

OpenStudy (anonymous):

sorry.. i'm not great in algeb, only in geom. xD

OpenStudy (anonymous):

this is geometry

OpenStudy (anonymous):

oh.,, let me see. xD

OpenStudy (anonymous):

As long as you got a side length and two angles or two side lengths and one angle, you can use the \[A.side/\sin(\angle.\alpha)=B.side/\sin(\angle.\beta)=C.side/\sin(\angle.\delta)\] And since sin(90deg) is always 1, this commen value will always be the length of the hypotenuse.

OpenStudy (anonymous):

but i only have one side and one angle measure

OpenStudy (anonymous):

cant i use cos or tan

OpenStudy (anonymous):

i know one angle because it is a right triangle. so one is 90. say one said is 7. the hypo side. couldnt i use tan

OpenStudy (anonymous):

If you only have those two value, I personaly can't see how it can be solved. If not a third side or angle we can get a value from, is there proportions?

OpenStudy (anonymous):

Is it a right sided or right angled triangle?